Product Pricing and Profit Analysis with Bottleneck Operations

Hercules Steel Company produces three grades of steel: high, good, and regular grade. Each of these products (grades) has high demand in the market, and Hercules is able to sell as much as it can produce of all three. The furnace operation is a bottleneck in the process and is running at 100% of capacity. Hercules wants to improve steel operation profitability. The variable conversion cost is $15 per process hour. The fixed cost is $200,000. In addition, the cost analyst was able to determine the following information about the three products:

High
Grade
Good
Grade
Regular
Grade
Budgeted units produced 5,000 5,000 5,000
Total process hours per unit 12 11 10
Furnace hours per unit 4 3 2.5
Unit selling price $280 $270 $250
Direct materials cost per unit $90 $84 $80

The furnace operation is part of the total process for each of these three products. Thus, for example, 4.0 of the 12.0 hours required to process High Grade steel are associated with the furnace.

Required:

1. Determine the unit contribution margin for each product.

Contribution Margin
Per Unit
High Grade $
Good Grade $
Regular Grade $

2. Provide an analysis to determine the relative product profitability, assuming that the furnace is a bottleneck. Round your answers to two decimal places.

Contribution Margin
Per furnace Hour
High Grade $
Good Grade $
Regular Grade $
Answer #1
a
High Grade Good Grade Regular Grade
Unit selling price 280 270 250
Less: Variable costs
Direct materials cost per unit 90 84 80
Variable conversion cost 180 165 150
Contribution Margin Per Unit 10 21 20
b
High Grade Good Grade Regular Grade
Contribution Margin Per Unit 10 21 20
Divide by Furnace hours per unit 4 3 2.5
Contribution Margin Per furnace Hour 2.50 7.00 8.00
Workings:
High Grade Good Grade Regular Grade
Variable conversion cost =12*15 =11*15 =10*15